Two simple clutch disks of equal mass 6.3 kg are initially separate. They also have equal radii of R=0.55 m. One of the disks is accelerated to 5.7 rad/s in time Δt = 1.8 s. They are then brought in contact and both start to sping together. Calculate the angular velocity of the two disks together.


1
Expert's answer
2022-05-07T14:40:30-0400

"M_1=M_2,"

"\omega_2=\frac{\omega_1}{\sqrt 2}=4~\frac{rad}s."